A system consists of two spheres, of mass m and 2m
Answer #1

This problem the system is formed by two objects to be rigid

Part 29)

the momentum is

I1 =m v

and the system is linked by the bar together two two body falls

I2 = (2m) v

we can see that for the second body momentum is

I2 = 2 I1

so the total momentum is

It =I1 +I2

It = 5 + 10

It = 15 Kg m/s

result correct B

Part 30)

The Earth accelerates the two bodies the acceleration of gravity, but it creates a different torque on each

The center of mass

Xcm =( m X1 +2m (L-X1) )/3m

Xcm = X1 + 2 L – 2 X1 )/3

Xcm = 2L/3 – X1 /3

whereby the center of mass is to 2L / 3 and L smaller mass / 3 of the heavier mass

whereby the torque

τ = F x

τ = (mg) 2L/3 – (2mg) L/3 =0

consequently the torque is canceled by which the bar does not turn in his fall

result correct C